OpenStudy (anonymous):

Could someone help Me Make And Equation For This? I'll Give a Metal , And ill Fan <3 Please!! ^.^ Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation

OpenStudy (anonymous):

let the amount of water be x then

OpenStudy (anonymous):



OpenStudy (anonymous):

where did you get 3?

OpenStudy (anonymous):

Actually the amount of apple juice is 3 times of quantity of water!

OpenStudy (anonymous):

@jigglypuff314 can you help? im confused

jigglypuff314 (jigglypuff314):

he means that 25 times 3 = 75 25% = 1L so 75% = 3L so you get 3L + 1L = 100%

OpenStudy (anonymous):

i too did the same! @Courtneydyrdek98

OpenStudy (anonymous):

so if i put that id get it right?

OpenStudy (anonymous):

let the total volume be x litres theres 75% apple juice (75/100)*x 0.75x total volume=0.75x+1 x=0.75x+1

OpenStudy (anonymous):

thank you you think you could help me with the other part?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (anonymous):

0.25x=1 x=4 theres 0.75x litres apple juice present in the container 0.75*x

OpenStudy (anonymous):

thank you dear (: